Mica travel constant speed 3 hours. Then slowed speed for 11 miles. The entire trip took 6 hours and covered 375 miles. What was the speed on the turnpike?
Answer by ankor@dixie-net.com(22740)

Mica travel constant speed 3 hours.
Then slowed his speed by 11 mph. (Changed the wording here to make sense.)
The entire trip took 6 hours and covered 375 miles.
What was the speed on the turnpike?
:
Let s = the speed on the turn-pike
then
3s = his distance at turn-pike speed
the total time was 6 hr, therefore:
3(s-11) = distance at the slower speed
:
Write a distance equation; dist = speed * time
3s + 3(s-11) = 375
3S + 3S - 33 = 375
6S = 375 + 33
6S = 408
S = 408/6
S = 68 mph on the turn-pike